A certain experiment conducted was normally distributed. The mean value is 50 and the standard deviation is 4. If the total population is 1000 (discrete variable), calculate the probability for getting an x-value greater than 45


1
Expert's answer
2021-11-24T03:43:05-0500

Let X be a random value from the given population, then X ~ "N(50,4^2)=50+4N(0,1)"

"P(X>45)=P(50+4N(0,1)>45)=P(N(0,1)>-1.25)=0.89435"

There is 0.89435 probability of getting x-value greater than 45